OpenStudy (anonymous):

A large emerald with a mass of 812.04 grams was recently discovered in a mine. If the density of the emerald is 2.76grams over centimeters cubed, what is the volume? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point.

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

@phi @satellite73 @AravindG

OpenStudy (anonymous):

@Compassionate @iambatman @thomaster

OpenStudy (phi):

write down what you know \[ 2.76 \frac{\text{g}}{\text{cc}} \] 812.04 g notice if we multiply by a factor K that has units cc/g we would have \[ 812.04\ \cancel{g} \cdot K \frac{cc}{\cancel{g}} = K\cdot 812.04\]

OpenStudy (phi):

* \(K \cdot 812.04 \) cc so you need to find a factor that has units of cc/g that you can multiply by. any idea what that factor is ?

OpenStudy (anonymous):

@phi my internet cut sorry

OpenStudy (anonymous):

@phi 2.76

OpenStudy (phi):

first, in these kinds of problem *always* include the units we know 2.76 g/cc but we want cc/g any ideas what to do ?

OpenStudy (anonymous):

Uh so maybe 812.04?

OpenStudy (phi):

812.04 g we want cc/g

OpenStudy (anonymous):

sorry bit lost

OpenStudy (phi):

notice this \[ \frac{1}{2.76 \frac{g}{cc} }\] if we multiply top and bottom by cc we get \[ \frac{1\cdot cc}{2.76 \frac{g}{\cancel{cc}} \cdot \cancel{cc}} = \frac{1}{2.76} \frac{cc}{g} \]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ganeshie8 my internet keeps turning off so pls dont leave if i disconect

OpenStudy (anonymous):

ill reconect ASAP